# 3DGAA: Realistic and Robust 3D Gaussian-based Adversarial Attack for Autonomous Driving

### v2

@@ 1 Introduction
Camera-based object detection systems serve as a cornerstone of autonomous driving [2, 1, 37], enabling essential tasks such as obstacle avoidance and lane navigation. However, increasing researches show that these models are highly vulnerable to adversarial examples. In physical world, subtle, structured perturbations often induce misclassifications [10, 4, 11, 30, 46, 28]. This raises serious concerns about the reliability of perception modules in safety-critical scenarios. Thus, systematically exploring such vulnerabilities is essential to understand and improve the real world safety of autonomous driving systems [3, 7, 38, 45, 15, 42, 40, 39, 32, 33, 16, 26].
@@ 2 Related Work
To improve generalization, recent works leverage 3D object representations and render adversarial textures onto full object meshes [45, 42, 40, 39, 32]. In particular, FCA [39] generates full-surface camouflage with physical constraints. Environment-aware enhancements [32, 33] further boost robustness, and recent works such as TT3D [16] and PGA [26] explore high-dimensional texture search spaces.
@@ 3.5 Physical Augmentation Module
3.6 Selective Dimension Optimization
To support flexible adversarial sample generation under physical constraints, we introduce a manual dimension selection strategy that decouples geometric and appearance parameters in the 3DGS space. Specifically, users can define a subset of dimensions $K\subset\{1,\dots,14\}$ to be optimized, allowing certain parameters to remain fixed based on deployment requirements. This design enables tailored perturbation control across different real-world scenarios. For instance, one may constrain geometry when targeting printable objects or preserve opacity to ensure sensor transparency. The strategy also provides practical benefits in training efficiency and transferability, as further validated in Section 4.5. Implementation details and additional visualizations are deferred to Appendix A.4 for clarity.
@@ 3.6 Selective Dimension Optimization
5:     $I\leftarrow R(\theta)$triangleright Render current 3DGS
6:     $\tilde{I}\leftarrow\mathcal{T}(I)$triangleright Apply physical augmentation
@@
9:     $\theta\leftarrow\theta-\text{Mask}_{\mathcal{K}}(\delta)$triangleright Update selected dimensions only
@@ 4.1 Experimental Setup
Baseline Methods. We compare 3DGAA against representative open-source 2D and 3D physical attack baselines, including CAMOU[45], UPC[15], DAS[40], FCA[39], DTA[32], ACTIVE[33] and TT3D[16]. For a fair comparison, all methods are adapted to the same synthetic and real-world setups with standard calibration and consistent rendering settings.
Generation Pipeline and Camera Setup. For digital rendering, we use a differentiable 3D Gaussian renderer following [34]. All adversarial objects are rendered under multi-view camera settings, with 12 evenly spaced azimuth angles and 3 distances (3m, 5m, 10m), simulating realistic surveillance conditions. For physical-world evaluation, we deploy textured miniature models under varied lighting conditions and capture images using a Realme GT Neo5 SE camera mounted on a motorized rig with fixed focal length in a proportionally reduced distance.
@@
Evaluation Metrics. We report attack effectiveness using the Log Confidence Reduction (LCR), which measures the proportion of images where the detector fails to assign the correct label to the adversarial object. LCR is defined as:
@@
Unlike absolute confidence drop, LCR normalizes adversarial effectiveness across different baselines, ensuring a fairer comparison. We also report mean Average Precision with 50% threshold (mAP@0.5) to reflect detection performance decline. To quantify physical realism, we adopt image similarity metrics including LPIPS [44], SSIM [41], and PSNR, computed between rendered adversarial objects and their original clean versions.
Implementation Details. All experiments are conducted on a single NVIDIA RTX 4090 GPU. We set the number of epochs in 50, the attack training finished in 1 minute (see Appendix C.3). The two loss terms in our objective–adversarial loss $\mathcal{L}_{\text{adv}}$ and shape consistency loss $\mathcal{L}_{\text{shape}}$–are dynamically weighted during optimization. Specifically, we compute the relative loss magnitudes and normalize them to balance gradients: the weight for $\mathcal{L}_{\text{adv}}$ is inversely proportional to its scaled loss value, while the weight for $\mathcal{L}_{\text{shape}}$ is proportional to its raw value. This dynamic scheme eliminates the need for manual tuning of fixed loss weights.
Table 1: Benchmark Comparison of 3DGAA and existing baseline adversarial attack methods. Our method 3DGAA achieves the best performance across almost all metrics, demonstrating superior physical realism and adversarial robustness.
Method Vanilla CAMOU[45] UPC[15] DAS[40] FCA[39] DTA[32] ACT.[33] TT3D[16] 3DGAA
LPIPS $\downarrow$ 0.0000 0.6210 0.6161 0.5979 0.5629 0.6142 0.6105 0.5433 0.5373
SSIM $\uparrow$ 1.0000 0.0212 0.0224 0.0132 0.0123 0.0198 0.0193 0.0122 0.0128
PSNR $\uparrow$ $\infty$ 0.1829 0.1749 0.4590 0.2987 0.1804 0.1814 0.3876 0.4951
LCR $\uparrow$ 0.0000 0.3128 0.2095 0.5366 1.2810 0.8784 1.3778 1.1202 3.5628
mAP $\downarrow$ 87.21% 70.21% 75.42% 60.12% 35.89% 47.44% 33.56% 40.12% 7.38%
@@ 4.2 Comprehensive Comparison with Benchmark Physical Attacks
Physical Realism. As shown in Table 1, 3DGAA achieves the lowest LPIPS score (0.5373), indicating minimal perceptual distortion and high-fidelity texture appearance. While our SSIM score is marginally lower than that of UPC, this is primarily due to the inherent approximation in 3DGS-based geometry rather than the adversarial manipulation. In particular, 3DGAA achieves the highest PSNR value among attack methods, reflecting superior preservation of pixel-level fidelity during optimization. These results demonstrate the benefit of our shape loss and physical filtering module in maintaining object realism.
Attack Strength. 3DGAA significantly outperforms all baselines in terms of adversarial robustness in Table 1. It achieves an LCR of 3.5628, far exceeding the second-best method (ACTIVE at 1.3778), and reduces the detector’s mAP from 87.21% to just 7.38%. This indicates that our method not only fools the detector more often, but also leads to lower residual confidence, effectively suppressing recognition. These results validate the effectiveness of simultaneously optimizing both geometric and appearance parameters in the 3DGS space, enabling stronger and more transferable adversarial effects.
Viewpoint Robustness. We evaluate attack consistency by rendering adversarial objects from multiple viewpoints. As shown in Figure 3, the original object is consistently detected (top), while the 3DGAA version evades or misleads detection from all views (bottom). We also measure LCR across 12 azimuth angles and 3 distances. Figure 4 shows higher success rates at side views and close range, with performance slightly declining under front views or long distances, validating the multi-view training strategy and 3DGAA’s robustness to viewpoint variations.
Together, these results confirm that 3DGAA achieves strong attack success while preserving physical realism, and remains effective under diverse viewpoints, which are key properties for physically deployable adversarial objects.
Figure 3: Multi-Angle Adversarial Attack Effectiveness: Top row shows original samples with successful detection (green boxes) across different viewpoints. Bottom row demonstrates 3DGAA-generated adversarial textures evading detection (black boxes) or false detection (red boxes) at corresponding angles. The consistent performance across viewing perspectives highlights 3DGAA robustness to viewpoint variations.
@@ 4.3 Effectiveness of the Physical Filtering Module
Table 3: Ablation study of the shape loss component. $\mathcal{L}_{adv}+\mathcal{L}_{shape}$ achieves the best physical realism.
@@ 4.7 Physical Realization
To verify the real-world effectiveness of 3DGAA, we conduct physical deployment experiments using miniature 3D-printed vehicles under four lighting conditions: indoor daylight (ID), indoor night (IN), outdoor daylight (OD), and outdoor night (ON). As illustrated in Figure 7, we compare three texture variants: (1) vanilla (non-adversarial), (2) 3DGAA without physical augmentation, and (3) 3DGAA with augmentation. The vanilla object is reliably detected across all conditions. In contrast, 3DGAA w/o Aug. only slightly reduces detection confidence, particularly under lighting. The full 3DGAA w/ Aug. achieves consistent detection failures or misclassifications across all four scenarios. These results demonstrate the critical role of the physical augmentation module in bridging the sim-to-real gap.
We further report AP@0.5 across four detectors in Table 6. 3DGAA significantly reduces AP under real-world imaging from 84.17% to as low as 17.92%, surpassing all baseline physical attacks. This validates the full pipeline’s ability to generate physically robust adversarial objects deployable in uncontrolled physical environments.
@@ 5 Conclusion
A detailed discussion of security concerns and potential societal risks and mitigation strategies is included in Appendix C.4.
@@ A.1 Dynamic Loss Weighting Strategy
Let ¯$\bar{\mathcal{L}}_{\text{adv}}$ and ¯$\bar{\mathcal{L}}_{\text{shape}}$ denote the average adversarial loss and shape loss at a given iteration. We first scale the adversarial loss for numerical stability:
^$\hat{\mathcal{L}}_{\text{adv}}=\bar{\mathcal{L}}_{\text{adv}}/\gamma,$ (A.1)
@@ A.4 Selective Dimension Optimization
Geometry-Selected Mode ($\mathcal{K}_{g}$): This mode activates all 10 geometry-related dimensions, including position ($x$, $y$, $z$), scale ($s_{x}$, $s_{y}$, $s_{z}$), and quaternion rotation ($q_{w}$, $q_{x}$, $q_{y}$, $q_{z}$):
@@
Appearance-Selected Mode ($\mathcal{K}_{a}$): This mode only optimizes the 4 appearance-related parameters—color ($c_{r}$, $c_{g}$, $c_{b}$) and opacity ($\alpha$):
@@ B Supplementary of Experiments
B.1 Dataset Configuration
Due to the 3DGS generator $\mathcal{F}_{3DGS}$ was pretrained on synthetic scenes. Our framework requires only inference-time multiview inputs. We collect the images from Carla[8], covering: 20 vehicle models with varying geometries, 5 weather conditions (clear, rain, fog, etc.), Multi-scale observations and Full spherical viewpoints (24 azimuth/elevation angles).
B.2 Implementation Details
@@
B.3 Selective Dimension Optimization
@@
B.4 Object Generalization
A critical aspect of adversarial attacks is their ability to generalize across different object categories. To evaluate the generalization capability of 3DGAA, we apply our adversarial optimization framework to multiple vehicle types, including sedans, SUVs, trucks, and buses. As shown in Fig. B.3, the results demonstrate that 3DGAA consistently reduces the detection confidence of different vehicle types, confirming its robustness and transferability across diverse 3D object geometries.
These findings indicate that the adversarial perturbations learned by 3DGAA effectively adapt to various object structures, ensuring broad applicability in real-world scenarios. Moreover, the ability to generalize across object categories strengthens the practicality of 3DGAA in autonomous driving security research, where adversarial robustness must be considered for a wide range of vehicle shapes and sizes.
B.5 Multi-Angle Visualization of 3DGAA Results
@@ B.6 Human Perception
3DGAA* 7.1 3DGAA 7.9
3DGAA*: Appearance-Selected mode.
@@
We analyze the collected ratings by computing the average realism score per condition and compare different adversarial and vanilla textures. The results presented in Table B.1 demonstrate the superior realism of our proposed 3DGAA method compared to existing adversarial attack approaches. Notably, 3DGAA achieves a significantly higher realism score of 7.9, outperforming prior methods such as ACTIVE (6.0), UPC (6.4), and DTA (3.2), indicating that adversarial perturbations generated by 3DGAA are less perceptible to human observers. The appearance-selected variant, 3DGAA* (7.1), further validates the effectiveness of our approach, demonstrating that the texture modifications in 3DGAA contribute to improved realism while maintaining adversarial effectiveness. Compared to real vehicle textures (Vanilla: 9.6), our method achieves the closest resemblance, confirming that our Physical Filtering Module and Physical Augmentation Module successfully reduce unrealistic artifacts. Furthermore, conventional adversarial texture-based methods, such as CAMOU (2.9), DAS (3.9), and FCA (2.3), exhibit significantly lower realism scores, primarily due to their reliance on large-scale texture distortions that lack physical consistency. The extreme case of Random (1.1) highlights that naive perturbations are easily identifiable as artificial, further reinforcing the necessity of structured optimization in adversarial texture generation. These results demonstrate that 3DGAA achieves a superior balance between adversarial robustness and realism, making it more applicable to real-world scenarios where physical reality is crucial.

### v3

@@ 1 Introduction
Camera-based object detection systems serve as a cornerstone of autonomous driving [2, 1, 38], enabling essential tasks such as obstacle avoidance and lane navigation. However, increasing researches show that these models are highly vulnerable to adversarial examples. In physical world, subtle, structured perturbations often induce misclassifications [10, 4, 11, 31, 48, 28]. This raises serious concerns about the reliability of perception modules in safety-critical scenarios. Thus, systematically exploring such vulnerabilities is essential to understand and improve the real world safety of autonomous driving systems [3, 7, 39, 46, 15, 43, 41, 40, 33, 34, 16, 26, 47, 29].
@@ 2 Related Work
To improve generalization, recent works leverage 3D object representations and render adversarial textures onto full object meshes [46, 43, 41, 40, 33, 47, 29]. In particular, FCA [40] generates full-surface camouflage with physical constraints. Environment-aware enhancements [33, 34] further boost robustness, and recent works such as TT3D [16] and PGA [26] explore high-dimensional texture search spaces.
@@ 3.5 Physical Augmentation Module
3.6 Dimension Selection for Physical Realization
Real-world deployment imposes a non-deformability constraint on vehicles. To explicitly accommodate this, we expose a dimension-selection switch in 3DGAA that partitions the 3DGS parameters into $\theta_{\text{app}}=\{\text{rgb},\text{opacity}\}$ (appearance) and $\theta_{\text{geo}}=\{\text{pos},\text{scale},\text{rot}\}$ (geometry), and instantiate three modes:
Appearance-only $\text{3DGAA}^{a}$. We freeze $\theta_{\text{geo}}$ and optimize only $\theta_{\text{app}}$. In this setting, the resulting adversarial object can be fabricated by printable wraps/films without any geometric alteration (see Physical Realization). Technically, this mode achieves the best perceptual fidelity among attack methods (Table 1: lowest LPIPS and highest PSNR.
Geometry-only $\text{3DGAA}^{g}$. We fix $\theta_{\text{app}}$ and optimize $\theta_{\text{geo}}$ under tight displacement bounds implicitly encoded by $L_{\text{shape}}$. This ablates the contribution of geometric cues and keeps competitive realism (Table 1) while exposing how pose/scale perturbations contribute to detector failures.
Full (3DGAA). We jointly optimize $\theta_{\text{app}}\cup\theta_{\text{geo}}$ with $L_{\text{shape}}$. This configuration provides the strongest attack effect (highest LCR, lowest mAP in Table 1) at a small realism cost relative to $\text{3DGAA}^{a}$).
The switchable design aligns optimization with physical constraints: $\text{3DGAA}^{a}$ directly supports geometry-free, printable deployment on vehicles; $\text{3DGAA}^{g}$ isolates geometric sensitivity; and 3DGAA realizes the upper bound of attack efficacy. In practice, $\text{3DGAA}^{g}$ converges faster and with lower memory due to frozen appearance, while 3DGAA achieves the best robustness. Moreover, $\text{3DGAA}^{a}$ serves as the methodological precursor to our fabrication pipeline (see Physical Realization) and the comprehensive comparisons in Exp. 4.2 and 4.7.
@@ 3.6 Dimension Selection for Physical Realization
5:  $I\leftarrow R(\theta)$ $\triangleright$ Render current 3DGS
6:  $\tilde{I}\leftarrow\mathcal{T}(I)$ $\triangleright$ Apply physical augmentation
@@
9:  $\theta\leftarrow\theta-\text{Mask}_{\mathcal{K}}(\delta)$ $\triangleright$ Update selected dimensions only
@@ 4.1 Experimental Setup
Baseline Methods. We compare 3DGAA against representative open-source 2D and 3D physical attack baselines, including CAMOU [46], UPC [15], DAS [41], FCA [40], DTA [33], ACTIVE [34], TT3D [16], RAUCA [47] and PACG [29]. For a fair comparison, all methods are adapted to the same synthetic and real-world setups with standard calibration and consistent rendering settings.
Object & Invariance. All physical variants use the same rigid miniature shell; geometry is unchanged and only textures differ (Sec. 4.7). Textures used for fabrication are produced by the appearance-only pipeline $\text{3DGAA}^{a}$ (Sec. 3.6).
Synthetic Scene Selection. We adopt CARLA [8] sampling with fixed seeds to collect the images, covering 20 vehicle models with varying geometries, 5 weather conditions, multi-scale observations and full spherical viewpoints (24 azimuth/elevation angles).
@@
Evaluation Metrics. We report attack effectiveness using the Log Confidence Reduction (LCR), which measures the proportion of images where the detector fails to assign the correct label to the adversarial object:
@@
Unlike absolute confidence drop, LCR normalizes adversarial effectiveness across different baselines, ensuring a fairer comparison. We also report mean Average Precision with 50% threshold (mAP@0.5) to reflect detection performance decline. To quantify physical realism, we adopt image similarity metrics including LPIPS [45], SSIM [42], and PSNR, computed between rendered adversarial objects and their clean counterparts. All metrics are computed under the same view and lighting protocol as in Sec. 4.2 to ensure fairness.
Implementation Details. All experiments run on a single NVIDIA RTX 4090 with fixed driver/CUDA/PyTorch versions and fixed seeds. We use early stopping under a shared rule; minutes-level wall-clock runtime profiling (including 3DGS generation, adversarial training, and I/O/misc) and peak memory under this protocol are summarized in App. C.3. The two loss terms in our objective–adversarial loss $\mathcal{L}_{\text{adv}}$ and shape-consistency loss $\mathcal{L}_{\text{shape}}$–are dynamically weighted by normalizing relative magnitudes to balance gradients: the weight for $\mathcal{L}_{\text{adv}}$ is inversely proportional to its scaled value, while the weight for $\mathcal{L}_{\text{shape}}$ is proportional to its raw value. This dynamic scheme eliminates manual tuning of fixed loss weights.
Generation Pipeline and Camera Setup. Digital rendering: we use a differentiable 3D Gaussian renderer following [35]. All adversarial objects are rendered under multi-view settings with 12 evenly spaced azimuth angles and 3 distances (3 m, 5 m, 10 m), aligned with the protocol in Sec. 4.2. Physical fabrication: textures are UV-unwrapped and printed on matte wraps to reduce glare, then applied to the same rigid shell (no body modification); materials, printing ppi, and alignment cues are summarized in App. B.1. Physical images are captured with a Realme GT Neo5 SE; camera/exposure specifics and layout sketches are provided in App. B.1.
Table 1: Benchmark Comparison of 3DGAA against existing physical adversarial attacks. Besides the full model (3DGAA), we report two dimension-selection variants: $\text{3DGAA}^{a}$ (appearance-only optimization) and $\text{3DGAA}^{g}$ (geometry-only optimization). Across metrics, our approach consistently delivers superior physical realism and adversarial robustness.
Method Vanilla CAM.[46] UPC[15] DAS[41] FCA[40] DTA[33] ACT.[34] TT3D[16] RAU.[47] PACG[29] $\textbf{3DGAA}^{a}$ $\textbf{3DGAA}^{g}$ 3DGAA
LPIPS $\downarrow$ 0.0000 0.6210 0.6161 0.5979 0.5629 0.6142 0.6105 0.5433 0.6056 0.6152 0.5218 0.5423 0.5373
SSIM $\uparrow$ 1.0000 0.0212 0.0224 0.0132 0.0123 0.0198 0.0193 0.0122 0.0178 0.0204 0.0123 0.0193 0.0128
PSNR $\uparrow$ $\infty$ 0.1829 0.1749 0.4590 0.2987 0.1804 0.1814 0.3876 0.2572 0.1956 0.5480 0.4741 0.4951
LCR $\uparrow$ 0.0000 0.3128 0.2095 0.5366 1.2810 0.8784 1.3778 1.1202 1.4705 1.0299 2.6766 3.4410 3.5628
mAP $\downarrow$ 87.21% 70.21% 75.42% 60.12% 35.89% 47.44% 33.56% 40.12% 31.47% 42.71% 13.64% 8.03% 7.38%
@@ 4.2 Comprehensive Comparison with Benchmark Physical Attacks
Physical Realism. As shown in Table 1, 3DGAAa attains the lowest LPIPS (0.5218), with 3DGAA (0.5373) and 3DGAAg (0.5423) following as second and third best, respectively. While our SSIM is marginally lower than UPC, this largely reflects the inherent approximation of 3DGS geometry rather than adversarial manipulation. Importantly, 3DGAAa achieves the highest PSNR (0.5480) among attack methods, with 3DGAA second (0.4951) and 3DGAAg third (0.4741), indicating superior pixel-level fidelity. These results highlight the benefit of our shape-consistency loss and physical filtering in preserving realism, and show that appearance-only optimization can further tighten perceptual fidelity.
Attack Strength. 3DGAA significantly outperforms all baselines in adversarial robustness (Table 1). It achieves an LCR of 3.5628, exceeding the strongest baseline RAUCA (1.4705), and reduces mAP from 87.21% to 7.38%. Both variants also surpass all baselines: 3DGAAg reaches 3.4410 LCR / 8.03% mAP and 3DGAAa 2.6766 LCR / 13.64% mAP. This confirms that optimizing both geometry and appearance yields the strongest effect, while our selective dimension optimization (Method 3.6) still delivers SOTA degradation.
Viewpoint Robustness. We measure LCR across 12 azimuth angles and 3 distances. Figure 3 shows higher success rates at side views and close range, with performance slightly declining under front views or long distances, validating the multi-view training strategy and 3DGAA’s robustness to viewpoint variations.
Together, these results confirm that 3DGAA achieves strong attack success while preserving physical realism, and remains effective under diverse viewpoints. Moreover, the strong results of 3DGAAa verify a purely appearance-based and geometry-free deployment path, which is critical for physically deployable adversarial objects.
@@ 4.3 Effectiveness of the Physical Filtering Module
Table 3: Ablation study of the shape loss component. $\mathcal{L}_{adv}+\mathcal{L}_{shape}$ achieves better physical realism.
@@ 4.7 Physical Realization
To verify real-world effectiveness under the non-deformability constraint, we deploy adversarial textures on the same miniature vehicle shell so that geometry remains unchanged across all variants. Textures are obtained from our pipeline $\text{3DGAA}^{a}$ (Sec. 3.6) and transferred to printable wraps. Practically, this constitutes a geometry-preserving, print-and-apply workflow: textures from $\text{3DGAA}^{a}$ are UV-unwrapped and printed on matte wraps applied to the same rigid shell (no body modification); the procedure scales to full-size vehicles by proportionally adjusting wrap size/ppi and view-distance while keeping the capture protocol fixed.
We evaluate four lighting conditions: indoor daylight (ID), indoor night (IN), outdoor daylight (OD), and outdoor night (ON), and compare three texture variants (Fig. 6): (1) vanilla (non-adversarial), (2) 3DGAA w/o Aug. (no physical augmentation), and (3) 3DGAA w/ Aug. The vanilla object is reliably detected in all scenes. Without augmentation, adversarial textures mildly reduce confidence yet show sensitivity to illumination. In contrast, 3DGAA w/ Aug. consistently induces missed detections or misclassifications across ID/IN/OD/ON, indicating that our augmentation–modeling exposure/lighting/partial occlusion during optimization–effectively bridges the sim-to-real gap.
Table 6 quantifies detector performance (AP@0.5) on four architectures. With identical geometry and capture protocol, 3DGAA reduces AP from 84.17-96.25% down to 17.92-35.83%, surpassing all physical baselines. Together with Sec. 3.6, these results establish a geometry-preserving, appearance-only fabrication path for vehicles while retaining strong attack efficacy under diverse real-world lighting.
@@ 5 Conclusion
A discussion of security concerns and potential societal risks and mitigation strategies is included in App. C.5.
@@ A.1 Dynamic Loss Weighting Strategy
Let $\bar{\mathcal{L}}_{\text{adv}}$ and $\bar{\mathcal{L}}_{\text{shape}}$ denote the average adversarial loss and shape loss at a given iteration. We first scale the adversarial loss for numerical stability:
$\hat{\mathcal{L}}_{\text{adv}}=\bar{\mathcal{L}}_{\text{adv}}/\gamma,$ (A.1)
@@ A.4 Selective Dimension Optimization
Geometry-only Mode ($\mathcal{K}_{g}$): This mode activates all 10 geometry-related dimensions, including position ($x$, $y$, $z$), scale ($s_{x}$, $s_{y}$, $s_{z}$), and quaternion rotation ($q_{w}$, $q_{x}$, $q_{y}$, $q_{z}$):
@@
Appearance-only Mode ($\mathcal{K}_{a}$): This mode only optimizes the 4 appearance-related parameters—color ($c_{r}$, $c_{g}$, $c_{b}$) and opacity ($\alpha$):
@@ B Supplementary of Experiments
B.1 Implementation Details
@@
B.2 Selective Dimension Optimization
@@
B.3 Object Generalization
A critical aspect of physical attacks is their ability to generalize across objects. We evaluate 3DGAA on multiple vehicle types (sedan, SUV, truck, bus). As shown in Fig. B.3, our framework consistently reduces detector confidence across these categories, indicating robustness under diverse 3D geometries.
Cross-vehicle adaptation routes. Because 3DGS couples appearance to a scene-specific Gaussian set (and implicitly to geometry/UV layout), cross-vehicle use is achieved via two lightweight routes: (i) Initialization transfer: use an already optimized source model as initialization on the target model and run a few optimization steps to align views (Sec. 3.6); (ii) UV reparameterization: map the source texture onto the target via UV alignment (piecewise or learned correspondence) and optionally refine with brief optimization. Both routes are compatible with the appearance-only setting $\text{3DGAA}^{a}$ (no body modification) and benefit from our minutes-level runtime (App. C.3), making per-model adaptation practical.
These findings suggest that 3DGAA’s adversarial perturbations adapt effectively across rigid vehicles by choosing appearance-only deployment; for cones/barriers, either appearance-only or simple geometry shells are feasible. Overall, 3DGAA trades a universal single-texture assumption for a controllable, geometry-aware procedure that remains broadly applicable while preserving physical deployability.
B.4 Multi-Angle Visualization of 3DGAA Results
@@ B.5 Human Perception
RAUCA 3.7 PACG 3.6
$\text{3DGAA}^{a}$ 7.1 3DGAA 7.9
$\text{3DGAA}^{a}$: Appearance-only mode.
@@
We analyze the collected ratings by computing the average realism score per condition and compare different adversarial and vanilla textures. The results presented in Table B.1 demonstrate the superior realism of our proposed 3DGAA method compared to existing adversarial attack approaches. Notably, 3DGAA achieves a significantly higher realism score of 7.9, outperforming prior methods such as ACTIVE (6.0), UPC (6.4), and RAUCA (3.7), indicating that adversarial perturbations generated by 3DGAA are less perceptible to human observers. The appearance-selected variant, $\text{3DGAA}^{a}$ (7.1), further validates the effectiveness of our approach, demonstrating that the texture modifications in 3DGAA contribute to improved realism while maintaining adversarial effectiveness. Compared to real vehicle textures (Vanilla: 9.6), our method achieves the closest resemblance, confirming that our Physical Filtering Module and Physical Augmentation Module successfully reduce unrealistic artifacts. Furthermore, conventional adversarial texture-based methods, such as CAMOU (2.9), DAS (3.9), and FCA (2.3), exhibit significantly lower realism scores, primarily due to their reliance on large-scale texture distortions that lack physical consistency. The extreme case of Random (1.1) highlights that naive perturbations are easily identifiable as artificial, further reinforcing the necessity of structured optimization in adversarial texture generation. These results demonstrate that 3DGAA achieves a superior balance between adversarial robustness and realism, making it more applicable to real-world scenarios where physical reality is crucial.